# ADPO: Anchored Direct Preference Optimization A Unified Framework from Pairwise to Listwise Preferences

### v3

@@ Abstract
Direct Preference Optimization (DPO) has become a standard for aligning models with human feedback, yet its reliance on hard, pairwise preferences makes it brittle to annotator noise and distribution shift. We propose Anchored Direct Preference Optimization (ADPO), a generalized framework that learns from soft, listwise supervision by anchoring policy updates to a reference model. Our key theoretical contribution is to show that this anchoring mechanism imposes an implicit trust region on the policy update, enforced by the softmax Fisher information metric. This provides a robust geometric interpretation for both fixed and dynamic anchor strategies.
@@ 1 Introduction
To address these limitations, we introduce Anchored Direct Preference Optimization (ADPO), a unified framework that generalizes preference learning to soft and listwise supervision. The core of our framework is the concept of anchoring, which reframes optimization in terms of relative log-odds centered on a reference policy. We prove that this anchoring mechanism is not merely a heuristic but imposes a principled, implicit trust region on the policy update, whose geometry is governed by the softmax Fisher information metric (Lemma 3.3). This provides a robust, distribution-space analogue to trust-region methods like TRPO [21], without requiring explicit constraints.
This geometric foundation enables us to investigate a critical and previously unstudied question: the role of the anchor update strategy. We provide the first systematic comparison of a fixed anchor (frozen throughout training) versus a dynamic anchor (periodically updated to track the current policy). Our central finding is a crucial task-dependent trade-off:
@@
Collectively, our work makes four key contributions: (1) we propose a unified framework for preference optimization that robustly handles soft and listwise data; (2) we establish the theoretical connection between anchoring and implicit, Fisher-metric trust regions; (3) we uncover the task-dependent trade-off between anchor strategies, a key practical insight; and (4) we provide extensive empirical evidence and clear guidance for selecting anchor strategies in different learning contexts, informing both online RLHF and offline distillation pipelines.
@@ 3 Anchored Direct Preference Optimization
3.1 Pairwise Soft-DPO
@@
Remark 3.1 (Scale non-identifiability).
Temperatures $\beta$ and $\beta_{r}$ only appear through ratio $\beta_{r}/\beta$ in the optimum. We set $\beta=\beta_{r}=1.0$ by default.
Proposition 3.2 (Special cases).
Soft-DPO recovers: (i) standard DPO when $q_{ij}\in\{0,1\}$ and $\Delta_{\text{ref}}=0$; (ii) non-anchored Bradley–Terry when $\Delta_{\text{ref}}=0$; (iii) reward-based soft preferences when $q_{ij}=\sigma(\beta_{r}(R_{i}-R_{j}))$.
Implication. This shows that ADPO is a strict generalization of standard DPO, and the anchoring term $(s-s^{\mathrm{ref}})$ reduces to unanchored log-odds when $\Delta_{\text{ref}}=0$. Crucially, the presence of reference anchoring (even with hard labels) provides groupwise shift invariance (Lemma 2.1), which standard DPO lacks.
Proof of (i).
For $q_{ij}=1$: $\ell_{ij}=\log(1+e^{\beta\Delta_{\theta}})-\beta\Delta_{\theta}=-\log\sigma(\beta\Delta_{\theta})$. For $q_{ij}=0$: $\ell_{ij}=\log(1+e^{\beta\Delta_{\theta}})=-\log\sigma(-\beta\Delta_{\theta})$. This matches standard DPO exactly. ∎
3.2 Listwise Soft-DPO

### v4

@@ 1 Introduction
To address these limitations, we introduce Anchored Direct Preference Optimization (ADPO), a unified framework that reveals a fundamental connection across seemingly disparate learning paradigms. Through a simple anchored projection formulation—minimizing $\mathrm{KL}(q\|\tilde{p}_{\theta})$ where $\tilde{p}_{\theta}$ is defined via anchored logits $(s-s^{\mathrm{ref}})/\tau$—we show that supervised fine-tuning, knowledge distillation, maximum-entropy RL policy updates, advantage-weighted methods (AWR/AWAC/MPO), and DPO all emerge as special cases through different choices of target distribution $q$, anchor policy $\pi_{\mathrm{ref}}$, and temperature $\tau$ (Section 3, Proposition 3.1). The anchoring structure is key: by centering the coordinate system at a reference policy before applying softmax, ADPO provides groupwise shift invariance and Fisher-metric regularization—properties absent in standard methods.
We prove that this anchoring mechanism provides a principled geometric framework: the softmax Fisher information metric governs the local curvature around the target distribution (Lemma 3.3), yielding an implicit trust region centered at the target (not the anchor itself). This provides a robust, distribution-space analogue to trust-region methods like TRPO [23], without requiring explicit constraints. Crucially, this geometric foundation holds regardless of anchor choice, enabling us to investigate the role of the anchor update strategy. To our knowledge, this is the first systematic comparison of a fixed anchor (frozen throughout training) versus a dynamic anchor (periodically updated to track the current policy) in the context of preference learning and policy optimization. Our central finding is a crucial task-dependent trade-off:
@@
Collectively, our work makes four key contributions: (1) we propose ADPO as a unified projection framework that subsumes major learning paradigms (SFT, KD, max-ent RL, DPO) and robustly handles soft and listwise data; (2) we establish the theoretical connection between anchoring and implicit, Fisher-metric trust regions; (3) we uncover the task-dependent trade-off between anchor strategies, a key practical insight; and (4) we provide extensive empirical evidence and clear guidance for selecting anchor strategies in different learning contexts, informing both online RLHF and offline distillation pipelines.
@@ 3 Anchored Direct Preference Optimization
Notation.
Throughout this section, we use the following notation:
  • $x$: context (prompt/state); $S_{x}=\{1,\ldots,|S_{x}|\}$: candidate set (actions/outputs)
  • $\pi_{\theta}(i\mid x)$: student policy probability for candidate $i$ given context $x$
  • $\pi_{\mathrm{ref}}(i\mid x)$: reference (anchor) policy, either fixed or dynamic
  • $s_{i}=\log\pi_{\theta}(i\mid x)$, $s_{i}^{\mathrm{ref}}=\log\pi_{\mathrm{ref}}(i\mid x)$: log-probabilities
  • $q(i\mid S_{x})$: target preference distribution over candidates (from teacher, rewards, or labels)
  • $u_{i}=(s_{i}-s_{i}^{\mathrm{ref}})/\tau$: anchored logits
  • $\tilde{p}_{\theta}(i\mid S_{x})=\mathrm{softmax}(u)$: anchored model distribution
Table 1: Temperature parameters and their roles across ADPO formulations.
Symbol Role Where used
$\beta_{r}$ Teacher/target inverse temperature Forms $q$ from rewards: $q\propto\exp(R/\beta_{r})$
$\beta$ Student inverse temperature (pairwise) Pairwise logistic: $\sigma(\beta(\Delta_{\theta}-\Delta_{\mathrm{ref}}))$
$\tau$ Anchoring temperature Defines anchored logits: $u=(s-s^{\mathrm{ref}})/\tau$ (listwise)
Remark (scale non-identifiability): In pairwise settings, temperatures $\beta$ and $\beta_{r}$ only appear through their ratio $\beta_{r}/\beta$ in the Bayes-optimal condition, so the statistical optimum is invariant to jointly rescaling them. We set $\beta=\beta_{r}=1.0$ by default. However, the optimization dynamics are affected: pairwise gradients scale as $\beta$ (since $\partial\ell/\partial\theta=\beta(\sigma(\beta(\Delta_{\theta}-\Delta_{\mathrm{ref}}))-q_{ij})(\nabla s_{i}-\nabla s_{j})$), while listwise gradients scale as $1/\tau$ (because $\partial\mathcal{L}/\partial s_{i}=\tfrac{1}{\tau}(\tilde{p}-q)$). Thus smaller $\beta$ slows pairwise updates; smaller $\tau$ tightens the trust region via curvature $\propto 1/\tau^{2}$ but increases gradient magnitudes.
Remark (temperature $\tau$ and regularization strength): In listwise settings with anchoring, $\tau$ controls the balance between the anchor and target via the closed-form solution $\pi_{\theta}\propto\pi_{\mathrm{ref}}\cdot q^{\tau}$ (Proposition 3.2). Intuitively:
  • As $\tau\to 0$: $q^{\tau}\to\mathbf{1}$ (uniform), so $\pi_{\theta}\to\pi_{\mathrm{ref}}$ (stronger regularization toward the anchor).
  • As $\tau\to\infty$: $q^{\tau}$ becomes increasingly peaked at $\operatorname*{arg\,max}q$, so $\pi_{\theta}$ tracks the target $q$ more aggressively.
Note that $\tilde{p}_{\theta}$ always matches $q$ at optimality regardless of $\tau$; the temperature affects how the underlying policy $\pi_{\theta}$ relates to the anchor. Smaller $\tau$ induces a tighter implicit trust region in the sense that large deviations $\log(\pi_{\theta}/\pi_{\mathrm{ref}})$ are more heavily penalized (since gradients scale as $1/\tau$).
3.1 ADPO as a Unified Anchored Projection
We cast preference learning as an anchored projection on the probability manifold: given target $q$ and anchor $\pi_{\mathrm{ref}}$, minimizing $\mathrm{KL}(q\|\tilde{p}_{\theta})$ aligns the student under the softmax Fisher geometry. This unified view encompasses supervised fine-tuning, knowledge distillation, maximum-entropy RL, and DPO as special cases.
Master formulation.
Given student $\pi_{\theta}$, anchor $\pi_{\mathrm{ref}}$, and temperature $\tau>0$, the anchored logits are
$u_{i}\;=\;\frac{s_{i}-s_{i}^{\mathrm{ref}}}{\tau},\qquad s_{i}=\log\pi_{\theta}(i\mid x),\ \ s_{i}^{\mathrm{ref}}=\log\pi_{\mathrm{ref}}(i\mid x).$ (3)
Note (sequence models): In autoregressive sequence models, we use length-normalized log-probabilities (token-average NLL) for $s$ and $s^{\mathrm{ref}}$ to avoid length bias. Under this setting, each candidate’s score is the per-token average, and these averaged scores enter the group-level softmax over candidates. Since length normalization amounts to an equal rescaling of all scores within a group, it preserves the softmax affine-invariance: the Fisher geometry and anchoring properties remain intact when applied to this sequence-level distribution over candidates (not the token-level distribution). Thus the theoretical results—groupwise shift invariance (Lemma 2.1), implicit trust region (Lemma 3.3), and closed-form projection (Proposition 3.2)—carry through in this setting. The anchored model distribution over a group $S_{x}$ is
$\tilde{p}_{\theta}(i\mid S_{x})\;=\;\frac{\exp(u_{i})}{\sum_{j\in S_{x}}\exp(u_{j})}\;=\;\mathrm{softmax}_{i}\Big(\tfrac{\log\pi_{\theta}(i\mid x)-\log\pi_{\mathrm{ref}}(i\mid x)}{\tau}\Big).$ (4)
Given a target preference distribution $q(\cdot\mid S_{x})$ (pairwise or listwise), ADPO minimizes
$\mathcal{L}_{\mathrm{ADPO}}\;=\;\mathbb{E}_{x,S_{x}}\!\left[\mathrm{KL}\!\big(q(\cdot\mid S_{x})\,\|\,\tilde{p}_{\theta}(\cdot\mid S_{x})\big)\right]\;=\;\mathbb{E}_{x,S_{x}}\left[-\sum_{i\in S_{x}}q(i|S_{x})\log\tilde{p}_{\theta}(i|S_{x})\right]+\mathrm{const}.$ (5)
This objective performs a projection onto the anchored probability manifold; its second-order expansion around $q$ yields the softmax Fisher metric, inducing an implicit trust region centered at $\pi_{\mathrm{ref}}$ (formalized in Section 3.4).
Relation to standard DPO.
DPO [1] also introduces a reference model, but it enters only through log-odds differences between two candidates:
$\Delta_{\theta,\mathrm{ref}}=\big(\log\pi_{\theta}-\log\pi_{\mathrm{ref}}\big)_{y_{i}}-\big(\log\pi_{\theta}-\log\pi_{\mathrm{ref}}\big)_{y_{j}}.$
In contrast, ADPO anchors each score $s_{i}-s_{i}^{\mathrm{ref}}$ individually before normalizing across the group to form $\tilde{p}_{\theta}$. When $|S_{x}|=2$ and $q$ is binary (or Bradley–Terry soft targets), Eq. (5) reduces to the logistic preference loss used in DPO. For general listwise $|S_{x}|>2$, ADPO retains distribution-level anchoring and inherits a Fisher-metric trust region around $\pi_{\mathrm{ref}}$.
Proposition 3.1 (ADPO unifies major learning paradigms).
Let $\mathcal{L}_{\mathrm{ADPO}}$ be defined in Eq. (5). Different choices of target $q$, anchor $\pi_{\mathrm{ref}}$, and temperature $\tau$ recover:
  1. 1.
    Supervised fine-tuning (SFT). If $q(\cdot\mid S_{x})$ equals the data label distribution (one-hot or soft), $\pi_{\mathrm{ref}}$ is uniform (constant logits that cancel in softmax), and $\tau=1$, then $\tilde{p}_{\theta}=\pi_{\theta}$ and $\mathcal{L}_{\mathrm{ADPO}}=\mathbb{E}[\mathrm{CE}(q,\pi_{\theta})]$, the standard cross-entropy loss. For $\tau\neq 1$, $\tilde{p}_{\theta}$ is a temperature-scaled version of $\pi_{\theta}$.
  2. 2.
[…]
  3. 3.
    Maximum-entropy RL / offline policy projection.
    • (Soft-optimal Boltzmann policy) The standard maximum-entropy RL objective yields the soft-optimal policy $\pi^{\star}(a|s)\propto\exp(Q(s,a)/\alpha)$ [19, 21, 20]. To match this target via ADPO’s anchored distribution $\tilde{p}_{\theta}$, we set $q(a|s)\propto\exp(Q(s,a)/\alpha)$ and use a uniform anchor $\pi_{\mathrm{ref}}$.
      Parametrization: ADPO constrains the anchored distribution $\tilde{p}(u)=\mathrm{softmax}((\log\pi_{\theta}-\log\pi_{\mathrm{ref}})/\tau)$ to equal $q$. Since there are equivalent parameterizations, we choose $\tau=1$ and $\pi_{\theta}(a|s)\propto\pi_{\mathrm{ref}}(a|s)\exp\!\big(Q(s,a)/\alpha\big)$ (with $\pi_{\mathrm{ref}}$ uniform, this reduces to $\pi_{\theta}\propto\exp(Q/\alpha)$). This directly recovers the standard soft-optimal policy $\pi^{\star}$ for use in sampling/evaluation, preserving the conventional Boltzmann temperature $1/\alpha$. Alternative parameterizations (e.g., $\tau=\alpha$ with $\pi_{\theta}\propto\exp(Q)$) yield the same $\tilde{p}$ but would shift the temperature into the anchoring scale, which we avoid for clarity.
    • (AWR/AWAC/MPO advantage-weighted projection) When $S_{x}$ is the full action space with overlapping support between $\pi_{\text{old}}$ and the target $q$, the anchor is $\pi_{\mathrm{ref}}=\pi_{\text{old}}$, and the target is $q(a|s)\propto\pi_{\text{old}}(a|s)\exp\!\big(A^{\pi_{\text{old}}}(s,a)/\beta\big)$ where $A^{\pi_{\text{old}}}=Q^{\pi_{\text{old}}}-V^{\pi_{\text{old}}}$ (shift-invariant to constant baselines), with $\tau=\beta$. The optimal anchored distribution is $\tilde{p}_{\theta}=q\propto\pi_{\text{old}}\exp(A/\beta)$, exactly matching the advantage-weighted target.
      Derivation of $\pi_{\theta}$ form: ADPO enforces $\tilde{p}(u)=q$ with $u=(\log\pi_{\theta}-\log\pi_{\text{old}})/\tau$ and $\tau=\beta$. Then
      $\tilde{p}(u)\ \propto\ \exp\!\Big(\frac{\log\pi_{\theta}-\log\pi_{\text{old}}}{\beta}\Big)\ \stackrel{{\scriptstyle!}}{{\propto}}\ \pi_{\text{old}}\,\exp\!\big(A/\beta\big),$
      which is equivalent to
      $\frac{\pi_{\theta}}{\pi_{\text{old}}}\ \propto\ \Big(\pi_{\text{old}}\,e^{A/\beta}\Big)^{\beta}\ \Rightarrow\ \pi_{\theta}\ \propto\ \pi_{\text{old}}^{\,1+\beta}\ e^{A}.$
      What is exactly recovered? ADPO matches the anchored distribution $\tilde{p}$ to the advantage-weighted target $q$; the induced policy shape $\pi_{\theta}\propto\pi_{\text{old}}^{\,1+\beta}\exp(A)$ is a consequence of the anchored projection (Prop. 3.2) and is not the classic MPO/AWAC M-step update (which yields $\pi_{\theta}\propto\pi_{\text{old}}\exp(A/\beta)$ or directly $\pi_{\theta}=q$ under sufficient capacity). Our equivalence is thus at the $\tilde{p}-$level rather than the $\pi_{\theta}$-level. Assumptions: (1) full action space $S_{x}$ is enumerable or covered by samples; (2) advantage weights $\exp(A/\beta)$ are normalizable; (3) for off-policy data $d^{\beta}$, we approximate by ignoring or clipping importance ratios (standard in AWR/AWAC [24, 25]), yielding weighted MLE $\max_{\theta}\mathbb{E}_{(s,a)\sim d^{\beta}}[\exp(A/\beta)\log\pi_{\theta}(a|s)]$. MPO’s E-step weights $w\propto\exp(Q/\eta)$ with dual variable $\eta$ (enforcing a KL constraint via Lagrangian duality) follow the same projection form [26], though MPO additionally solves for $\eta$ to satisfy a KL budget.
  4. 4.
    DPO (binary comparison form). For $|S_{x}|=2$ with candidates $\{i,j\}$ and Bradley–Terry soft targets $q(i\mid\{i,j\})=\sigma(\beta_{r}\Delta R)$ where $\Delta R=R_{i}-R_{j}$ is the reward difference, Eq. (5) is equivalent to DPO up to temperature scaling and reference shift. Specifically, the Bayes-optimal condition $\sigma(\beta(\Delta_{\theta}-\Delta_{\mathrm{ref}}))=q_{ij}$ yields $\Delta_{\theta}-\Delta_{\mathrm{ref}}=(\beta_{r}/\beta)\Delta R$. When $\beta=\beta_{r}$ and $\pi_{\mathrm{ref}}$ is uniform (so $\Delta_{\mathrm{ref}}=0$), this is exactly identical to the original DPO objective; otherwise, it differs by a temperature rescaling $\beta_{r}/\beta$ and an anchor-induced shift $\Delta_{\mathrm{ref}}$.
Table 2: ADPO as a unifying projection framework. Different choices of target $q$, anchor $\pi_{\mathrm{ref}}$, and temperature $\tau$ recover major learning paradigms. Key: Optimization matches the anchored distribution $\tilde{p}_{\theta}=\mathrm{softmax}((s-s^{\mathrm{ref}})/\tau)$ to the target $q$; the corresponding student policy $\pi_{\theta}$ satisfies $\pi_{\theta}\propto\pi_{\mathrm{ref}}\cdot q^{\tau}$ when the optimum is reachable (Prop. 3.2).
Task Target $q(\cdot\mid x)$ Anchor $\pi_{\mathrm{ref}}$ $\tau$ Recovers ($\tilde{p}_{\theta}$ / $\pi_{\theta}$)
SFT data (one-hot/soft) uniform 11 cross-entropy / same
KD (standard) $\mathrm{softmax}(z_{T}/T)$ uniform $T$ $T^{2}\,\mathrm{CE}\!\big(q_{T},\mathrm{softmax}(s/T)\big)$ / train at $T$, infer T=1T{=}1
KD (anchored) teacher $\pi_{T}$ $\pi_{T}$ any $\mathrm{KL}\!\big(\pi_{T}\|\mathrm{softmax}((s-s_{T})/\tau)\big)$ / trust region
[…]
DPO (binary) Bradley–Terry / hard uniform any original DPO (pairwise)
ADPO (listwise) Plackett–Luce / soft any (teacher/old) any this work
Proposition 3.2 (Closed-form optimal solution).
Suppose $S_{x}$ covers the full action/class space and $q(\cdot\mid S_{x})\succ 0$. (Discrete case) If the policy $\pi_{\theta}$ has sufficient expressivity (i.e., the parametrization can represent arbitrary log-odds over $S_{x}$), then the distribution-level optimum of $\operatorname*{arg\,min}_{\theta}\mathrm{KL}(q\|\tilde{p}_{\theta})$ satisfies
$\pi_{\theta}(i\mid x)\;\propto\;\pi_{\mathrm{ref}}(i\mid x)\cdot q(i\mid x)^{\tau}.$ (6)
(Continuous case) For continuous action spaces, assume $q$, $\pi_{\theta}$, and $\pi_{\mathrm{ref}}$ admit densities with respect to a common base measure (e.g., Lebesgue measure). Define the anchored distribution via the density ratio:
$\tilde{p}_{\theta}(a)\;\propto\;\exp\!\Big(\frac{\log\pi_{\theta}(a)-\log\pi_{\mathrm{ref}}(a)}{\tau}\Big)\;=\;\Big(\frac{\pi_{\theta}(a)}{\pi_{\mathrm{ref}}(a)}\Big)^{1/\tau}.$
Setting $\tilde{p}_{\theta}=q$ yields the closed-form solution
$\pi_{\theta}(a)\;\propto\;\pi_{\mathrm{ref}}(a)\cdot q(a)^{\tau},$
which requires: (i) overlapping support: $q(a)>0\Rightarrow\pi_{\mathrm{ref}}(a)>0$ almost everywhere; (ii) integrability: $\int\pi_{\mathrm{ref}}(a)\cdot q(a)^{\tau}\,da<\infty$ to ensure normalizability. Under these conditions, the solution is unique up to the normalization constant (equivalently, the Radon–Nikodym derivative $d\pi_{\theta}/d\mu$ is determined by the multiplicative form, where $\mu$ is the base measure).
Proof (discrete): $\tilde{p}_{\theta}=\mathrm{softmax}\big((s-s^{\mathrm{ref}})/\tau\big)=q$ if and only if $(s-s^{\mathrm{ref}})/\tau=\log q+c\mathbf{1}$ for some constant $c$ (reflecting the shift-invariance of softmax), which rearranges to $\pi_{\theta}\propto\pi_{\mathrm{ref}}\,q^{\tau}$. The constant $c$ is absorbed into the proportionality constant. ∎
Implication and reachability. Equation (6) strictly characterizes the "anchored projection" as a multiplicative fusion of reference and target when the optimum is reachable. In discrete finite-action settings with tabular or sufficiently expressive function approximation, this optimum is exactly attainable. However, in structured settings (e.g., autoregressive language models with shared parameters across tokens, or policy networks with limited capacity), the optimum may be unreachable, and gradient descent performs an approximate projection in the parametric family. The KD, MaxEnt, and AWR special cases in Proposition 3.1 all follow directly from (6).
Listwise subset constraints and convexity. When $S_{x}$ is only a subset (pairwise/listwise ranking), (6) holds within each group, but the global $\pi_{\theta}$ must simultaneously satisfy multiple group constraints. Crucially, $\mathcal{L}_{\mathrm{ADPO}}$ is convex with respect to the anchored logits $u=(s-s^{\mathrm{ref}})/\tau$ within each group (as shown in Theorem 3.6), but the optimization over shared parameters $\theta$ is typically non-convex. The group-wise closed-form solution provides a target, but there may not exist a parameter $\theta$ that simultaneously achieves $\tilde{p}_{\theta}(\cdot|S_{x})=q(\cdot|S_{x})$ for all groups (except in the tabular case or with unlimited expressivity). Thus, in practice, ADPO performs a parameterized fitting that minimizes the expected group-wise KL, balancing the constraints across all training groups.
By centering the coordinate system at $\pi_{\mathrm{ref}}$ before applying softmax, ADPO provides groupwise shift invariance (Lemma 2.1) and Fisher-metric regularization (Lemma 3.3)—properties absent in standard KD or unanchored DPO. Table 2 summarizes the unification.
Standing assumptions.
Throughout this work, we maintain the following conditions unless otherwise noted:
  1. (i)
    Support overlap (discrete): For discrete action/candidate spaces, we assume $q(i|x)>0\Rightarrow\pi_{\mathrm{ref}}(i|x)>0$ for all $i\in S_{x}$, ensuring that anchored logits $u_{i}=(s_{i}-s_{i}^{\mathrm{ref}})/\tau$ are well-defined.
  2. (ii)
    Absolute continuity (continuous): For continuous action spaces, we assume $q$, $\pi_{\theta}$, and $\pi_{\mathrm{ref}}$ admit densities with respect to a common base measure, with overlapping support in the same sense as (i).
  3. (iii)
    Local analysis vs. global bounds: Fisher-metric trust region results (Lemma 3.3, Corollary 3.4) hold locally around $u^{\star}$ where $\tilde{p}_{\theta}(u^{\star})=q$. For global strong convexity, we either work in a neighborhood of $u^{\star}$ or assume the optimization is constrained to a bounded domain where $\tilde{p}_{\theta}(u)_{i}\in[\varepsilon,1-\varepsilon]$ (Theorem 3.6).
  4. (iv)
    Sequence models: For autoregressive sequence generation, we use length-normalized (per-token average) log-probabilities $s$ and $s^{\mathrm{ref}}$ as group-level scores, and apply the softmax Fisher geometry at the sequence-level distribution over candidates.
  5. (v)
    Expressivity: Closed-form optimal solutions (Proposition 3.2) assume the policy parametrization $\pi_{\theta}$ has sufficient expressivity to represent the target multiplicative form $\pi_{\mathrm{ref}}\cdot q^{\tau}$. In practice (e.g., shared-parameter networks, autoregressive models), we perform an approximate projection in the parametric family.
3.2 Pairwise Instantiation
Notation for this subsection: $\beta$ = student inverse temperature, $\beta_{r}$ = teacher/reference inverse temperature, $q_{ij}$ = soft preference probability that $i\succ j$.
@@ 3.2 Pairwise Instantiation
3.3 Listwise Instantiation
Notation for this subsection: $\tau$ = anchor temperature, $\beta_{r}$ = teacher inverse temperature for forming $q$ from rewards, $\tilde{R}_{i}$ = observed rewards, $\hat{R}_{i}$ = transformed rewards.